1. Company A sold 100,000 shares in an initial public offering. The underwriter's explicit fees were $90,000. The offering price for the shares was $30, but immediately upon issue, the share price jumped to $43. What is the net proceeds this company receives from this offering? 2. If you are pessimistic about stock A and its market price is $100/share. You short sell 1,000 shares. Suppose the stock A price falls to $70/share. You close your position and what is your profit? 3. A bond has a total five years maturity and has been there for two years since issuance. What should be its price as of now?
Answer #1

1. Company A sold 100,000 shares in an initial public offering. The underwriter's explicit fees were $90,000. The offering price for the shares was $30, but immediately upon issue, the share price jumped to $43. What is the net proceeds this company receives from this offering?

Net proceeds = Nos. of shares sold x Offering price - underwriter's fees = 100,000 x 30 - 90,000 = $  2,910,000

2. If you are pessimistic about stock A and its market price is $100/share. You short sell 1,000 shares. Suppose the stock A price falls to $70/share. You close your position and what is your profit?

Your profit = (Short sell price - purchase price) x number of shares short sold = (100 - 70) x 1000 = $ 30,000

3. A bond has a total five years maturity and has been there for two years since issuance. What should be its price as of now?

It's price should be present value of all the future (balance) coupon payments and principal redemption calculated using current / prevailing interest rate as discount rate. In the absence of any data, it will not be possible to quantify the numbers and the price.